OpenStudy (anonymous):

if a team consist of 20 members in which 20 % are mathematician then how many member should they hire so that the ratio would be 10/11

OpenStudy (anonymous):

calculate how many members are mathematicians they have first

OpenStudy (anonymous):

(they have 4)

OpenStudy (tkhunny):

Try to remember that your answer needs to be a Whole Number. Mathematicians don't like it when you define them in a piecewise fashion.

OpenStudy (anonymous):

@offnenStudieren : Aftr that... ?

OpenStudy (anonymous):

\[\frac{ 4+x }{ 20+x } = \frac{ 10 }{ 11 }\]